OpenStudy (anonymous):

The first term of geometric progression is two thirds of the sum to infinity. Find the common ratio.

hartnn (hartnn):

1st term = a sum to infinity = a/(1-r) r = common ratio Given : a= (2/3) a/(1-r) can u find r from here ?

OpenStudy (anonymous):

\[S \infty = \frac{ a }{ 1-r }\]

OpenStudy (anonymous):

Could u please help me further?

hartnn (hartnn):

a= (2/3) a/(1-r) cancelling a. 1 = 2/3 /(1-r) 1-r = 2/3 r = 1-2/3 = ??

hartnn (hartnn):

r=1/3